Shift Summary for Day shift, Oct 30, 1999

Miguel Barrio (mbarrio@hep.uchicago.edu)
Sat, 30 Oct 1999 15:47:28 -0500

Shift Summary for Day shift, Oct 30, 1999

On Shift: Miguel Barrio

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

No beam this shift due to TeV studies.

Activities:

08:00 Walk-through. Nothing looks obviously wrong.
He bag 1 is deflated, pressure in the bag reads zero. Since
DaveN was futzing with them, this is probably alright. Will try
to contact DaveN later to make sure. All other He bags look
fine.
DC gas flow is a little low, but this again seems to be fine
according to previous shifts.

12:00 DaveN came by the control room, he's inflating He bag 1 back
again.

13:40 Went through the gas checklist, everything's fine.

14:55 Walk-through. Everything looks honki-dori.
He bag 1 is back in it's puffy shape, dunno about the gas
mixture inside, though, but DaveN is taking care of that.

=========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: 0

-- 
Miguel Barrio
Email:  mbarrio@hep.uchicago.edu
Office: HEP 319, 773-702-7484; fax X-1914
Pager:  630-722-6854